Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
ciardelli, dattilo, sposito - 17:25 Thursday 22 August 2019 (46746) Print this report
Installation of heating belts at NI-WI
We took advantage of the commissining shift of this morning to install the heating belts on the lower part of the NI and WI vacuum towers. They are needed for the implementation of the etalon control.
We worked from the Hymo platform in the lower part of the clean rooms.
For the time being we left th belts disconnected.
Furthermore, the electrical cables from the heating belts to the TCS room are installed, and for the time being they too are left disconnected at both ends.
Activity concluded a few minutes before the lock of 10:05 UTC.

Images attached to this report
Comments to this report:
dattilo, montanari - 18:47 Tuesday 10 September 2019 (46882) Print this report

During this morning maintenance we completed the installation of the HW for the etalon control:

- installation of the drivers for the heating belts (Kert 420 model, Figs 1-2).
- installation of the remaining cables (prepared by G. Sposito).
- labeling and connections of all the parts.

The Kert420  have been connected to the mains via the remotely controlled EnerGenie PDU "TCS1", port #3 for NI, and port #4 for WI (Fig3).

In agreement with A.Masserot, the Kert inputs have been  connected to the  DAC #40  (DAQ box #31),  channel #2 for NI, and channel #3 for WI  (Fig4) for remote voltage adjustment..

The load seen by the Kert is 2.2 Ohms  (heating belt + cable).

With Alain at phone acting on the DAC values, we checked the remote adjustment of the voltage delivered to the heating belts:

Vdac Vout kert
1 2.2
5 11
10 22

So, as it can be evinced, tha maximum power that can be delivered to each heating belt is 220W.  This is enough to increase the T of the tower lower part (and of the payload)  of a few degrees.

At the end, we left  both Kert switched on, and with 0 voltage at the output.

They are ready to be remotely controlled by the DAC. 

Images attached to this comment
dattilo, sposito, cavalieri - 12:43 Thursday 12 September 2019 (46905) Print this report

During the maintenance of Sept 3rd, we finalized the cabling of  the temperature sensors attached to the bottom part of NI and WI towers (Figs 1-2).

They are PT100 sensors and are connected to their conditioning electronics placed in the dedicated  Dbox  on the tower rack of each tower. The analog outputs are  then sent to the IMMS box of DAQ room. 

We also performed a few calibration  measurements using precise resistors, summary in Fig3, to extract the  offset and slope values.

This morning, I updated the IMMS server cfg file with the new 2 channels and the corresponding slope and offset.

The name are INF..NI_BOTTOM_TE1 and  INF..WI_BOTTOM_TE1  (plot in Fig4).

Images attached to this comment
brooks, dattilo, masserot - 17:05 Tuesday 21 January 2020 (48210) Print this report

This morning during the maintenance monitor cables were added to the output of the north and west heating belt power supplies to monitor the voltage sent to the actuators. The following channels are now available:

  • LSC_NI_HB_moni
  • LSC_WI_HB_moni

Also, the clipped voltages were measured by using the analog gauge on the front of the power supply. The north was clipped at 22 V (at 10 A) as expected. The West was clipped at 18 V instead of the expected 22 V. After the maintenance we then realized that the voltage was clipped by the PyALP process which was controlling the west etalon control (the north had already migrated to Acl control). It is anticpated that the clipping voltage will be 22 V when not clipped by PyALP.

Search Help